I am thinking of a 2 digit number. The sum of the digits is 8. If I reverse the 2 digits then the value of the new number is 18 more than the orginal number. What is the orginal number?
-Do not understand, please help.

Answer by stanbon(75887)

I am thinking of a 2 digit number. The sum of the digits is 8. If I reverse the 2 digits then the value of the new number is 18 more than the orginal number. What is the orginal number?
-------
Let the number be 10t+u (t is the tens digit; u is the units digit)
-----
Equation:
t + u = 8
10u+t = 10t+u + 18
---------------------------
Substitute for "t" and solve for "u":
10u + (8-u) = 10(8-u)+u + 18
-----
10u - 9(8-u) = u+18
10u -72+9u = u+18
18u = 90
u = 5 (that is the units digit)
---
Solve for "t":
t = 8-u
t = 8-5
t = 3
----
Original Number: 10t+u = 35
==============================
Cheers,
Stan H.
================